### v1

@@ Abstract
Fast and accurate simulation of soft tissue deformation is a critical factor for surgical robotics and medical training. In this paper, we introduce a novel physics-informed neural simulator that approximates soft tissue deformations in a realistic and real-time manner. Our framework integrates Kelvinlet-based priors into neural simulators, making it the first approach to leverage Kelvinlets for residual learning and regularization in data-driven soft tissue modeling. By incorporating large-scale Finite Element Method (FEM) simulations of both linear and nonlinear soft tissue responses, our method improves neural network predictions across diverse architectures, enhancing accuracy and physical consistency while maintaining low latency for real-time performance. We demonstrate the effectiveness of our approach by performing accurate surgical maneuvers that simulate the use of standard laparoscopic tissue grasping tools with high fidelity. These results establish Kelvinlet-augmented learning as a powerful and efficient strategy for real-time, physics-aware soft tissue simulation in surgical applications.
Keywords:
Neural Simulators Soft Tissue Deformation Kelvinlet
1 Introduction
@@
Soft tissue deformation modeling is a critical need in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation supports surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional soft tissue deformation modeling relies on physics-based methods like Finite Element (FE) models, which offer high accuracy but are computationally expensive for real-time use [27]. Alternative models, including mass-spring systems [17], mesh-free methods [2], and particle-based methods [14], approximate tissue behavior, but require substantial computation and struggle to achieve realistic modeling.
@@
Recent advances in deep learning have enabled efficient, data-driven approximations of soft tissue deformations. CNNs, GNNs, and Transformers [20] can learn complex deformation patterns from simulations or real-world data, enabling real-time tissue simulation. However, they face two key limitations: 1) poor generalization to unseen force distributions and boundary conditions, leading to unreliable extrapolation, and 2) the risk of producing physically implausible deformations, particularly when trained on limited datasets.
To overcome these limitations, physics-informed machine learning integrates fundamental physical laws into neural network training, offering a promising approach to improving accuracy and generalization. Physics-Informed Neural Networks (PINNs) [22] enforce differential equations via soft constraints, while hybrid methods like differentiable FEM solvers [25] and coarse-to-fine neural surrogates [12] aim to balance accuracy and efficiency. However, these approaches are computationally demanding, particularly for large soft-tissue meshes, and require careful consideration of numerical stability and differentiability.
@@
We propose a physics-informed neural network framework that leverages Kelvinlets [6] to improve the accuracy and physical consistency of soft tissue deformation modeling. As a closed-form Green’s function solution to the Navier-Cauchy equations, Kelvinlets provide an efficient and analytically tractable prior. In this paper, we introduce two hybrid approaches: 1) using Kelvinlets as a baseline and learning the residual (modeling error) via a neural network, and 2) incorporating Kelvinlets as a regularizer for a neural network predicting the full deformation field. We generated a finite element dataset of liver deformations, using a linear model without preload and a nonlinear model with gravity-induced preload. We demonstrate that our proposed Neural-Augmented Kelvinlet outperforms purely data-driven models in generalization, physical realism, and learning efficiency, enabling real-time realistic modeling of soft-tissue deformations with multiple graspers.
Figure 1: Visualization of our Neural Augmented Simulator with realistic textures.
Contributions. Our work makes the following key contributions [^1] Repository available at https://github.com/Anon92373/Neural-Kelvinlet.:
@@
2 Kelvinlets for Soft Tissue Deformation
2.1 Navier-Cauchy Equations and Green’s Function Solution
where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response. For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
$\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\left(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\right)\mathbf{f},$ (2)
where $r=\|\mathbf{x}-\mathbf{x}_{s}\|$ denotes the Euclidean distance from the source.
$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\right)\mathbf{u}_{s}.$ (3)
For multiple interaction points, the Kelvinlet model represents the total organ deformation as $\mathbf{u}(\mathbf{x})=\sum_{i}k_{i}\mathbf{u}_{\varepsilon,i}(\mathbf{x}),$ where the coefficients $k_{i}$ are optimized to satisfy displacement boundary conditions $\mathbf{u}(\mathbf{x}_{i})=\mathbf{u}_{i}$, via:
$\operatorname*{argmin}_{\mathbf{k}}(\sum_{i}\|\sum_{j}k_{j}\mathbf{u}_{\varepsilon,j}(\mathbf{x}_{i})-\mathbf{u}_{i}\|^{2})+\lambda\sum_{i}k_{i}^{2}.$ (4)
This ensures that the contributions of individual Kelvinlet functions collectively satisfy the prescribed displacements. The regularization $\lambda=0.001$ ensures a minimum energy solution. The total Kelvinlet deformation model, including the optimization, can be performed in less than 3 ms for two simultaneously engaged graspers using native GPU acceleration in pytorch on a mesh with $\sim 10\text{k}$ nodes.
2.2 Method 1: Kelvinlet-Based Residual Learning
[…]
$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})-\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\right],$ (5)
where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain. By focusing on learning only the correction beyond the Kelvinlet solution, this approach enhances efficiency and interpretability while leveraging a known analytical model as a strong prior.
2.3 Method 2: Kelvinlets as Regularization
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]+\lambda_{reg}\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\epsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right],$ (6)
where $\lambda_{\text{reg}}>0$ is the regularization coefficient. For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation. Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (which are much faster). We will describe the distributions $p$ and $q$ used in our experiments in the following section.
For both proposed methods, when multiple graspers interact with the tissue, we employ the multi-point interaction solution of Kelvinlet and provide all interaction points along with their corresponding displacements to the network.
@@ 3 Large-Scale FEM Data
Our neural network framework trains on a large FE simulation dataset, focused on lifting liver lobes to expose the gallbladder during laparoscopic cholecystectomy. We construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using ‘gmsh’ [9].
Interactions of the liver and gallbladder with surrounding organs are modeled by prescribing boundary conditions. Those boundary conditions restrict regions of the liver’s exposed surface that are considered in the far field from the interaction location, as well as the distal end of the cystic ducts, arteries, and veins. FE simulations are solved using the open source getFEM package [24].
[…]
@@
In addition, this equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model, where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [10]:
$\boldsymbol{W}=C_{01}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-1}{3}}-3\right)+C_{10}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-2}{3}}-3\right),$ (8)
@@
$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),\vspace{-.1in}$ (10)
where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which is a function of the size of $\Omega_{i}$. The signals $u_{s}$ are independently sampled from a uniform distribution in a cuboid $\mathcal{C}$ defined by: $x\in(x_{1},x_{2}),y\in(y_{1},y_{2}),z\in(z_{1},z_{2})$, with volume: $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1}).$ This approach ensures that node sampling is spatially structured according to surgical priors while allowing displacements to be isotropic in a bounded range.
For single grasping, ten regions $\Omega_{i}$ around the liver edge were selected, each with 250 samples. For multiple graspers, six pairwise combinations of five regions $\Omega_{i}$ were considered, with 30 samples per region. The displacement ranges are $x\in(x_{1},x_{2})=(-90,50)$ mm, $y\in(y_{1},y_{2})=(-90,50)$ mm, and $z\in(z_{1},z_{2})=(-60,90)$ mm, corresponding to the cube domain.
@@
For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{s}^{1},\mathbf{x}_{s}^{2})=p(\mathbf{x}_{s}^{1})p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$, where $p(\mathbf{x}_{s}^{1})$ follows Eq. (10). To ensure that $\mathbf{x}_{s}^{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{s}^{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$ and adjusting $g(\Omega_{i\neq k})$ based on surgical priors, and then sampling $\mathbf{x}_{s}^{2}$ from $p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$.

### v2

@@ Abstract
Accurate and efficient modeling of soft-tissue interactions is fundamental for advancing surgical simulation, surgical robotics, and model-based surgical automation. To achieve real-time latency, classical Finite Element Method (FEM) solvers are often replaced with neural approximations; however, naively training such models in a fully data-driven manner without incorporating physical priors frequently leads to poor generalization and physically implausible predictions. We present a novel physics-informed neural simulation framework that enables real-time prediction of soft-tissue deformations under complex single- and multi-grasper interactions. Our approach integrates Kelvinlet-based analytical priors with large-scale FEM data, capturing both linear and nonlinear tissue responses. This hybrid design improves predictive accuracy and physical plausibility across diverse neural architectures while maintaining the low-latency performance required for interactive applications. We validate our method on challenging surgical manipulation tasks involving standard laparoscopic grasping tools, demonstrating substantial improvements in deformation fidelity and temporal stability over existing baselines. These results establish Kelvinlet-augmented learning as a principled and computationally efficient paradigm for real-time, physics-aware soft-tissue simulation in surgical AI. Our code and data is available at: https://github.com/Anon92373/Neural-Kelvinlet.
@@
Soft tissue deformation modeling is essential in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation enables surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional approaches rely on physics-based methods grounded in continuum mechanics, such as the Finite Element Method (FEM), Boundary Element Method (BEM), and Tensor Mass Model (TMM). While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications [2, 34]. To reduce computational cost, approximate methods such as mass-spring systems [23], mesh-free methods [3], and particle-based methods [21] are frequently used. However, these approximations often sacrifice realism, creating a significant simulation-to-reality gap that limits their practicality in high-fidelity applications.
@@
Recent advances in deep learning and neural operators have enabled fully data-driven methods that approximate soft tissue deformations with a single forward pass of a learned model. Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers [26] have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation. Despite these advances, fully data-driven approaches face two fundamental limitations: (a) limited generalization to unseen force distributions and boundary conditions, which results in unreliable extrapolation, and (b) a tendency to produce physically inconsistent deformations, particularly when trained on datasets of limited size or diversity. These shortcomings highlight the need for hybrid approaches that combine the expressiveness of neural models with the robustness of physics-based priors.
To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models [17]. Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks [5]; and (c) learning biases through the design of loss functions that explicitly enforce physical laws. Despite this progress, achieving the necessary low latency for real-time simulation while maintaining a high-performing physics-informed model with appropriately chosen biases remains a challenging and largely open problem.
In this paper, we model soft tissue using a volumetric mesh and address the problem of predicting the displacement of the entire mesh given the observed displacements at a limited set of interaction points on its surface. This is analogous to determining how the entire soft tissue will deform when graspers engage multiple surface locations and are moved to specified target positions. While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM [27], these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency. To overcome these limitations, we propose a physics-informed neural framework that integrates analytical elastic priors to enhance both accuracy and physical plausibility.
We introduce a physics-informed neural framework that leverages Kelvinlets [8], i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior. Kelvinlets provide a fast, low-order approximation of elastic deformation that captures the fundamental physics of linear elasticity. We build on this prior through two complementary hybrid strategies: (a) residual learning: using Kelvinlets to generate a baseline displacement field and training a neural network to model the residual error; and (b) physics regularization: incorporating Kelvinlets as a soft constraint within the learning objective of a neural network that predicts the full deformation field. By explicitly coupling the expressiveness of neural networks with the efficiency and physical grounding of Kelvinlets, our approach achieves improved accuracy, generalization, and physical plausibility for soft tissue deformation modeling.
@@
To establish a unified testbed for evaluating our approach alongside prior methods, we generated a finite element dataset of liver deformations using both a linear model without preload and a nonlinear model with gravity-induced preload. We show that our proposed Neural-Augmented Kelvinlet framework surpasses purely data-driven models in generalization, physical realism, and learning efficiency, ultimately enabling real-time, high-fidelity modeling of soft tissue deformations involving multiple graspers.
Contributions. Our key contributions are:
@@
1 Related Work
Classical approaches based on the Finite Element Method (FEM) provide high-fidelity simulations of soft tissue deformation but are often computationally prohibitive for real-time surgical applications. To address this limitation, model reduction techniques such as Proper Orthogonal Decomposition (POD) have been proposed to accelerate FEM without sacrificing physical realism. For example, Lau et al. [20] employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use. Complementarily, Hu and Desai [14] characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
More recently, physics-informed neural networks (PINNs) have emerged as a data-efficient alternative that integrates governing equations such as elasticity or dynamics directly into the learning objective. Hu et al. [15] introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility. Extending this paradigm, Nguyen-Le et al. [24] propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data. These methods improve interpretability and generalization but often require complex training and remain constrained by solver latency or limited scalability.
In contrast, purely data-driven surrogate models have been proposed to directly regress deformation fields from control variables. Ke et al. [6] train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference. While fast and simple, such models lack physical constraints, which can reduce reliability and robustness in unseen conditions.
On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery [33]. PEGNN [32] extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios. Meanwhile, SeeSaw [9] employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder. Zhao et al. [12] introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion. These approaches show strong empirical performance but may suffer from limited interpretability or reliance on dense supervision.
Neural architectures such as U-Net [22], Set Attention Block (SAB) [19], and GraphGPS [29] offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data. While effective in capturing spatial patterns and long-range dependencies, these models are generally agnostic to the underlying physics of soft tissue behavior and typically lack inductive biases grounded in biomechanics, such as force equilibrium, boundary conditions, or constitutive material priors.
Our work builds upon these foundations by integrating physically motivated priors, specifically Kelvinlet-based elastic solutions, into neural architectures through both residual learning and regularization. By embedding closed-form Green’s function responses from linear elasticity into the training process, our method encourages physically consistent deformation predictions even in data-sparse regimes. This hybrid approach improves generalization, improves physical plausibility, and remains compatible with a range of backbone architectures, while preserving the low latency performance required for real-time surgical simulation and interactive applications.
@@ 2 Large-Scale FEM Data
Figure 1: Overview of the large-scale FEM data generation pipeline. A patient’s abdominal CT scan is first anatomically segmented (left), followed by the generation of a 3D linear tetrahedral mesh (middle). Diverse interactions are then applied to the mesh, and FEM simulations are performed to produce training data for ML models (right). The right panel shows the norms of the input (interactions) and the resulting output displacements (predicted global effects).
Neural network–based solutions require training on large-scale finite element (FE) simulation datasets. In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh [11]. As part of our contributions, we release this large-scale dataset; the details of its creation are provided below.
The bare area of the liver, where it contacts the inferior vena cava and diaphragm, and the distal ends of the ducts and veins are fixed (non-deformable). Tool interaction regions are selected to simulate the initial surgical step of lifting the liver lobes to expose the gallbladder and cystic ducts. These regions are identified by clustering nodes along the liver ridge, based on surgical video analysis and input from our surgical collaborator. FE simulations are performed using the open-source getFEM package [30], applying tool-imposed displacements at those predefined clustered regions. This pipeline is illustrated in Figure 1.
In our FE model, nodes with restricted motion or tool-imposed displacements are flagged. These boundary conditions are encoded as binary features in the neural network (1 for constrained or controlled, 0 otherwise), allowing the model to distinguish between free and constrained regions.
Assuming a quasi-static environment, our solutions satisfy equilibrium by ensuring that the total potential energy is not only stationary but is also a minimum for permissible displacements, and is given by: $\int_{\Omega}\frac{\boldsymbol{W}(\boldsymbol{E})}{\partial\boldsymbol{\boldsymbol{E}}}:\delta\boldsymbol{E}\,dV+\int_{\Omega}\rho\boldsymbol{g}\cdot\delta\mathbf{u}\,dV=0$, where $\boldsymbol{W}(\boldsymbol{E})$ is the strain energy density, $\Omega$ is the volume of the model, $\delta\mathbf{u}$ is the variation of the displacement field, $\delta\boldsymbol{E}$ is the variation of the Green-Lagrange strain tensor, $\rho$ is the density of the material (1000 $kg/m^{3}$), and $\boldsymbol{g}$ is the acceleration due to gravity that points downward along the y-axis (9.81 $m/s^{2}$). For large displacements, the Green-Lagrange strain tensor $\boldsymbol{E}$ depends on the unknown displacement field $\mathbf{u}$ and is defined as:
@@
This equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model [31], where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [13]:
$\boldsymbol{W}=C_{01}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-1/3}-3\right)\\ +C_{10}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-2/3}-3\right),$ (2)
@@
$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),$ (4)
where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which depends on the size of $\Omega_{i}$. The displacements $\mathbf{u}_{s}\in\mathbb{R}^{3}$ are independently sampled from a uniform distribution over a cuboid domain $\mathcal{C}\subset\mathbb{R}^{3}$, defined by:
$x\in(x_{1},x_{2}),\quad y\in(y_{1},y_{2}),\quad z\in(z_{1},z_{2}),$
with total volume $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1})$. This ensures that node sampling is spatially structured according to surgical priors, while displacements remain isotropic within a bounded range.
For single grasping, ten regions $\Omega_{i}$ around the liver edge were selected, each with 250 samples. For multiple graspers, six pairwise combinations of five regions $\Omega_{i}$ were considered, with 30 samples per region. The displacement bounds are $x\in(-90,50)$ $m$, $y\in(-90,50)$ mm, and $z\in(-60,90)$ $m$.
@@
For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{1},\mathbf{x}_{2})=p(\mathbf{x}_{1})\,p(\mathbf{x}_{2}\mid\mathbf{x}_{1}),$ where $p(\mathbf{x}_{1})$ follows Eq. (4). To ensure that $\mathbf{x}_{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$, followed by renormalizing the remaining weights $g(\Omega_{i})$ for $i\neq k$ based on surgical priors. The conditional distribution $p(\mathbf{x}_{2}\mid\mathbf{x}_{1})$ is then defined using these updated region weights.